OpenStudy (anonymous):

If a phone company charges $0.45 for the first minute and $0.40 for each additional minute for a long-distance call, how much will a 33 -minute long-distance call cost?

OpenStudy (anonymous):

Okay, this is easy :) 33 minutes first minute is $0.45 32 minutes are at a rate of $0.40 per minute. 32 x .40 = 12 $12 + .45 = 12.45 $12.45 is the answer :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i guess it was wrong :S

OpenStudy (anonymous):

It was? Wth.

OpenStudy (anonymous):

If a phone company charges $0.50 for the first minute and $0.35 for each additional minute for a long-distance call, how much will a 32 -minute long-distance call cost? now they gave me this one

OpenStudy (anonymous):

0,45+(32*0,40)= 0,45*12,8=13,25 $

OpenStudy (anonymous):

Thanks Kate :/

OpenStudy (anonymous):

okay thank you kate